OpenStudy (anonymous):

This is a math based chemistry problem i anyone can help me plzz A sample of gaseous carbon dioxide was collected by the displacement of water at 228C and 736 torr. The volume under these conditions was measured to be 25.76 mL. How many moles of carbon dioxide are in this sample? How many millimoles of carbon dioxide are in this sample?

OpenStudy (anonymous):

pv=nrt

OpenStudy (anonymous):

iif anyone knows how to do it plz help me

OpenStudy (whpalmer4):

is there really any other kind of chemistry question? :-) okay, first, you'll want to convert that temperature to be in degrees K, not C. Do you know how?

OpenStudy (whpalmer4):

after you have the temperature converted, solve the ideal gas law for the value of \(n\) in terms of P, V, R, T and plug in the values. Be sure to do any unit conversion you need, and use a value of R that is compatible with your units.

OpenStudy (anonymous):

@whpalmer4 I solved it later but forgot to close the question I got my aswer as p.37 X10^-4 mol of co2

OpenStudy (anonymous):

or covertign it to millimoles the answer is .937 is that what u get?? I just want someone to double check!

OpenStudy (whpalmer4):

hmm, I get something else, can you show me your work?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

pv=nrt n=pv/rt p= 736/760 = .968 atm n= (.968)(.02444)/ (.0821)(295) n= 9.37X10 ^-4 mol

OpenStudy (whpalmer4):

you didn't convert the temperature to degrees K and your volume is a bit different than stated in the problem

OpenStudy (anonymous):

p=.968 23.44 ml into liters .02344 r=.0821 t= 22+273= 295

OpenStudy (anonymous):

so sorry that's a typo the one inn my problem is 23.44 mL

OpenStudy (whpalmer4):

The volume under these conditions was measured to be 25.76 mL A sample of gaseous carbon dioxide was collected by the displacement of water at 228C and 736 torr.

OpenStudy (anonymous):

and 22C sorry I copy and pasted it that's what I get.

OpenStudy (anonymous):

A sample of gaseous carbon dioxide was collected by the displacement of water at 22C and 736 torr. The volume under these conditions was measured to be 23.44 mL. How many moles of carbon dioxide are in this sample? How many millimoles of carbon dioxide are in this sample?

OpenStudy (anonymous):

okay but I am always use to converting C to K so that's what I did for thsi questiontoo

OpenStudy (whpalmer4):

ok, now the answers look a bit more similar :-)

OpenStudy (anonymous):

so id my answer right? so sorry about that typo u proabably had to redo it. sorry.

OpenStudy (whpalmer4):

I don't see anything wrong with it...

OpenStudy (anonymous):

tysm

OpenStudy (anonymous):

okay i have one more question but it's all chemistry

OpenStudy (whpalmer4):

ok

OpenStudy (anonymous):

define the "standard state of gases". If gases are usually not at "standard state", why does the phrase exist?

OpenStudy (whpalmer4):

you'll have to rummage around in your textbook or elsewhere for that, sorry

OpenStudy (anonymous):

yeah alredy did that idk where that stuff is

OpenStudy (whpalmer4):

wikipedia says The standard state for a gas is the hypothetical state it would have as a pure substance obeying the ideal gas equation at standard pressure (105 Pa, or 1 bar). No real gas has perfectly ideal behaviour, but this definition of the standard state allows corrections for non-ideality to be made consistently for all the different gases.

OpenStudy (anonymous):

thanks